A change purse contains an equal number of pennies, nickels, and dimes. The total value of the coins in $1.44. How many coins of each type does the purse contain?
Answer by richwmiller(17219)

p+5*n + 10*d = 144
p=n
n=d
p+5*p + 10*p = 144
16p=144
p=9
9p 9n 9d
9+45+90=144